OpenStudy (anonymous):

Express the series below in summation notation for the specified number of terms. 8. 6 + 11 + 16 + 21 + 26 + 31 + 36

ganeshie8 (ganeshie8):

start by finding out the \(n^{th}\) term

OpenStudy (anonymous):

\[\sum_{n=5}^{7} 6+n\]

ganeshie8 (ganeshie8):

nope

OpenStudy (anonymous):

why would it be wrong

ganeshie8 (ganeshie8):

is the given sequence arithmetic or geometric ?

OpenStudy (anonymous):

arithmetic

ganeshie8 (ganeshie8):

whats the formula for \(n^{th }\) term ?

ganeshie8 (ganeshie8):

6 + 11 + 16 + 21 + 26 + 31 + 36 first term, \(a = 6\) common difference, \(d = 5\) so, \(n^{th}\) term = \(a + (n-1)d = ?\)

OpenStudy (anonymous):

\[\sum_{n=1}^{7} a+(n-1)d\]

OpenStudy (anonymous):

so it would be like this

ganeshie8 (ganeshie8):

yes, plugin \(a\) and \(d\) values

ganeshie8 (ganeshie8):

\(\large \sum \limits_{n=1}^7 a+ (n-1)d\) \(\large \sum \limits_{n=1}^7 6+ (n-1)*5\)

ganeshie8 (ganeshie8):

simplify